Write an equation for the line in point/slope form and slope/intercept form that has the given condition.

6. Passes through (3,2) and is parallel to 2x-y=4


7. Passes through (-1,-1) and is perpendicular to y=5/2x+3
1
Expert's answer
2013-01-29T10:52:52-0500
The equation of the line is:
- in point/slope form y-y1=m(x-x1), where (x1,y1) - point on this line and m -
slope of this line;
- in slope/intercept form y=mx+b, where m - slope of this line and (0,b) -
coordinates of y-intercept
6. For this case: if two lines are parallel then their slopes are the same
=> m=2 and (x1,y1) = (3,2)
The equation for this line in point/slope form is y-2=2(x-3)
To get slope/intercept equation we have to transform the point/slope
equation to form y=mx+b:
y-2=2(x-3)
y=2x-6+2
y=2x-4
7. Similarly to previous case we have:
- in point/slope form: y+1=-2/5(x+1), 'cause for two perpendicular lines
m1=-1/m2, where m1 and m2 are the slopes of its lines;
- in slope/intercept form: y=-(2/5)x-2/5-1 y=-(2/5)x-7/5.